OpenStudy (anonymous):

what is the equation of the line parallel to -4x+2y=16 and passes through (9,-1) in slope intercept form?

OpenStudy (anonymous):

first get the slope.

OpenStudy (anonymous):

so get it into the form y=mx+b where m = slope so, y=8+2x and the slope is 8

OpenStudy (anonymous):

now you need a line that has the same slope but passes through (9,-1)

OpenStudy (anonymous):

start with point slope form (y-(-1))=8(x-9) and convert to slope intercept form. make sense?